What would be the displacement value of a drug, diclofenac if 20 suppositories containing 300 mg each of diclofenac weigh 43 g? One suppository made from only cocoa butter weighs 2 g.

0.16

1.1

0.5

2

1.16
Answer #1

Total paracetamol required for 20 suppositories 300*20= 6000 mg = 6 gms

total wt of 20 suppositories. 43 gms

weight of base required for 20 suppositories 2*20= 40 gms

displacement value of pcm. 6/3= 2